a wave travels in the x-direction its displacement y at point x in time t is y=0.1 sin (5x-6t) Its wavelength is? (a) π/5 (b) 2π/5 (c) π/2 (d) 5π/6 (e) 5π/3


1
Expert's answer
2022-02-23T08:32:54-0500

Given:

"y=0.1\sin (5x-6t)"


"y=A\sin \left(\frac{2\pi}{\lambda}x-\omega t\right)"

The wavelength

"\lambda=\frac{2\pi}{5}\:\rm m"

Answer: (b) 2π/5